# Directional excitability in Hilbert spaces

### v1

@@ I Introduction
To illustrate possible control applications of the proposed model, we test it as a directional spiking controller for a two-dimensional navigation task, in which case the spiking Hilbert space can be taken to be the two-dimensional Euclidean space. This toy example provides a simple sensorimotor testbed to explore the performance of neuromorphic controllers before their applications to more complex robotics situations. Several neuromorphic methodologies are already being developed for navigation tasks [12, 13], and the controller proposed here is akin the one considered in [14, 15]. The proposed model can be seen as a mid-point between the finite-dimensional, two-option (left or right) model used in [14] and the infinite-dimensional, infinite option (all possible planar directions) model used in [15]. The proposed model is finite-dimensional, yet capable of making spiking decision in any direction in the plane.
@@ II System definition and input-output characterization
$x_{s}$$\lVert\mathbf{x}\rVert$ $\frac{1}{\tau_{s}s+1}$ $g(\lVert\cdot{}\rVert)$ $\mathbf{u}(t)$$\mathbf{x}(t)$$x_{s}(t)$(1a)(1b)
Figure 1: Block diagram of the system (1). The red block at the top stands for gradient system (1a) and displays the bifurcation diagram of (4) with respect to $x_{s}$, where the full (resp. dotted) red lines are a continuum of stable (resp. unstable) equilibria. The two blue blocks at the bottom represent the slow dynamic (1b) modulating (1a).
@@
[…]
@@
System (1) consists of a gradient system (1a) modulated by the slow one-dimensional dynamic (1b). It is assumed to possess a solution on the maximum interval of existence $\mathbb{R}_{\geq 0}$ for all initial condition $(\mathbf{x}_{0},x_{s,0})\in\mathcal{H}\times\mathbb{R}_{\geq 0}$. The block diagram of (1) is presented in Figure 1.
@@
  • If $\underline{x}_{s}<x_{s}<\overline{x}_{s}$, then (4) has three stable equilibria at $r=-r^{*},0,r^{*}$. Furthermore, $\frac{\partial r^{*}}{\partial x_{s}}<0$.
@@
We have the following two theorems that characterizes the behavior of (4). Their proof can be found in the Appendix.
@@
Theorem 2 (Spiking direction dynamics)
@@ III Type I and II excitability
  • The $\lVert\mathbf{x}\rVert$-nullcline is the bifurcation diagram of (9a) with respect to $x_{s}$, as portrayed for $\mathbf{u}=0$ in the red block at the top of Figure 1.
@@ III-A Type I
  • For a constant input $\tilde{u}\in[0,\underline{\tilde{u}})$, the system acts as a simple integrator of the input near its resting equilibrium. A single spike can still be induced by a sufficiently strong transient input. The phase-plane of this behavior is portrayed at the top of Figure 4.
  • For a constant input $\tilde{u}\in(\underline{\tilde{u}},\overline{\tilde{u}})$, the system exhibits a tonic spiking activity characterized by an asymptotically stable limit cycle whose frequency grows with the input strength. The phase-plane of this behavior is portrayed at the middle of Figure 4.
  • For a constant input $\tilde{u}\in(\overline{\tilde{u}},+\infty)$, the system exhibits damped oscillations around the new globally asymptotically stable equilibrium. The phase-plane of this behavior is portrayed at the bottom of Figure 4.
Type I excitability is obtained with $\beta_{1}=3,\beta_{2}=1.5,\beta_{3}=1.5$. Using the nullcline analysis of (9), one can estimate that:
@@ III-B Type II
  • For a constant input $\tilde{u}\in[0,\underline{\tilde{u}})$, the system is globally asymptotically stable and acts as a simple resonator of the input near its resting equilibrium. A single spike can still be induced by a transient input which is either sufficiently large, or resonating with the system. The phase-plane of the near-to-equilibrium oscillations is portrayed at the top of Figure 5.
  • The cases $\tilde{u}\in(\underline{\tilde{u}},\overline{\tilde{u}})$ and $\tilde{u}\in(\overline{\tilde{u}},+\infty)$ are qualitatively similar to their Type I counterparts. The phase-planes of these behaviors are respectively portrayed at the middle and at the bottom of Figure 5.
@@ IV Application to $2d$ navigation with sparse actuation
with $1_{\lVert\mathbf{u}\rVert\geq S}\triangleq 1$ if $\lVert\mathbf{u}\rVert\geq S$ and $01_{\lVert\mathbf{u}\rVert\geq S}=$ 0 if $\lVert\mathbf{u}\rVert<S$. To foster sparse actuation and good sensitivity to environmental inputs, the controller dynamics are assumed to follow the generalized excitability model (1), as follows:
@@ IV-B Responsiveness to a changing environment
$\mathbf{v}=k_{1}\frac{\mathbf{z}^{*}-\mathbf{z}}{\varepsilon+d(\mathbf{z}^{*},\mathbf{z})}+k_{2}\sum_{i=0}^{n_{o}}\frac{\mathbf{z}-\mathbf{z}^{o_{i}}}{d(\mathbf{z},\mathbf{z}^{o_{i}})^{2}}$ (15)

### v2

@@ I Introduction
To illustrate possible control applications of the proposed model, we test it as a directional spiking controller for a two-dimensional navigation task, in which case the spiking Hilbert space can be taken to be the two-dimensional Euclidean space. This toy example provides a simple sensorimotor testbed to explore the performance of neuromorphic controllers before their applications to more complex robotics situations. Several neuromorphic methodologies are already being developed for navigation tasks [12, 13], and the controller proposed here is akin the one considered in [14, 15]. The proposed model can be seen as a mid-point between the finite-dimensional, two-option (left or right) model used in [14] and the infinite-dimensional, infinite option (all possible planar directions) model used in [15]. The proposed model is finite-dimensional, yet capable of making spiking decision in any direction of the plane.
@@ II System definition and input-output characterization
$_{s}𝐱\lVert\mathbf{x}\rVert$ $\frac{1}{\tau_{s}s+1}$ $g(\lVert\cdot{}\rVert)$ $\mathbf{u}(t)$$\mathbf{x}(t)$$x_{s}(t)$(1a)(1b)
Figure 1: Block diagram of the system (1). The red block at the top stands for gradient system (1a) and displays the bifurcation diagram of (4) with respect to $x_{s}$, where the full (resp. dotted) red lines are a continuum of asymptotically stable (resp. unstable) equilibria. The two blue blocks at the bottom represent the slow dynamic (1b) modulating (1a).
@@
  • The state $\mathbf{x}(t)$ and the input $\mathbf{u}(t)$ at each instant $t$ belong to a $\mathbb{R}$-Hilbert space $\left(\mathcal{H},\langle\cdot{},\cdot{}\rangle\right)$, with $\lVert\cdot{}\rVert$ the norm induced by the scalar product $\langle\cdot{},\cdot{}\rangle$.
@@
System (1) consists of a gradient system (1a) modulated by the slow one-dimensional dynamic (1b). It is assumed to possess a solution on the maximum interval of existence $\mathbb{R}_{\geq 0}$ for all initial condition $(\mathbf{x}_{0},x_{s,0})\in\mathcal{H}\times\mathbb{R}_{\geq 0}$. This can be demonstrated using the standard assumptions of the Cauchy-Lipschitz theorem [16]. The block diagram of (1) is presented in Figure 1.
Two key assumptions are imposed on the system. The first one is purely technical and serves only in the proof of Lemma 1. The second one is a necessary (though not sufficient) condition for the system to be excitable.
@@
  • If $\underline{x}_{s}<x_{s}<\overline{x}_{s}$, then (4) has three locally asymptotically stable equilibria at $r=-r^{*},0,r^{*}$. Furthermore, $\frac{\partial r^{*}}{\partial x_{s}}<0$.
@@
We have the following two theorems that characterize the behavior of (4). Their proof can be found in the Appendix.
@@
Theorem 2 (Direction dynamics)
@@ III Type I and II excitability
  • The $\lVert\mathbf{x}\rVert$-nullcline is the bifurcation diagram of (5a) (here (9a)) with respect to $x_{s}$, as portrayed for $\mathbf{u}=0$ in the red block at the top of Figure 1.
@@ III-A Type I
[…]
Here, Type I excitability is obtained with $\beta_{1}=3,\beta_{2}=1.5,\beta_{3}=1.5$. Using the nullcline analysis of (9), one can estimate that:
@@ III-B Type II
[…]
@@ IV Application to $2d$ navigation with sparse actuation
with $1_{\lVert\mathbf{u}\rVert\geq S}\triangleq 1$ if $\lVert\mathbf{u}\rVert\geq S$ and $1_{\lVert\mathbf{u}\rVert\geq S}=0$ if $\lVert\mathbf{u}\rVert<S$. To foster sparse actuation and good sensitivity to environmental inputs, the controller dynamic is assumed to follow the generalized excitability model (1), as follows:
@@ IV-B Responsiveness to a changing environment
$\mathbf{v}=k_{1}\frac{\mathbf{z}^{*}-\mathbf{z}}{\varepsilon+d(\mathbf{z}^{*},\mathbf{z})}+k_{2}\sum_{i=1}^{n_{o}}\frac{\mathbf{z}-\mathbf{z}^{o_{i}}}{d(\mathbf{z},\mathbf{z}^{o_{i}})^{2}}$ (15)